# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ V Analysis on Error bound
In this section, the error bound of the contribution function in the proposed EmSHAP is discussed. Theorem 1 indicates that the expectation of the mean absolute deviation (MAD) between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ has an upper bound of $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$, where $K$ is the number of Monte Carlo sampling, $\epsilon_{2}$ is a small positive number related to the absolute probability density ratio between the estimated conditional density and the true data-generating density. For comparison, Theorems 2 and 3 presents the error bounds of two competitive methods, i.e., KernelSHAP [12] and VAEAC [17].
@@
The parameters of the estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ are unbiased estimates of the parameters of the true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for any small positive $\epsilon_{1}$, the probability density ratio can be described as follows,
@@
Assumption 1 holds as EBM can be flexibly parameterized for arbitrary probability distributions [42, 43]. More specifically, the proposed GRU coupled energy-based model consists of an energy network and a GRU network. The energy network is an MLP with a skip connection and the GRU network is a type of gated recurrent network. According to Theorem 2.5 in Ref.[44], with weighted path norm uniformly bounded, by setting appropriate network width and depth, MLP with skip connection can approximate any function in Barron space [^2] Barron space is a function space that consists of functions which can be approximated by smooth combinations of simple basis functions.. On the other hand, Theorem 3.2 in Ref.[45] shows the universal approximation ability of GRU.
Before analyzing the upper error bound of EmSHAP, we first demonstrate the asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$, followed by the derivation of their error bounds.
V-A Asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$
@@
(Consistency) From the weak law of large numbers, when the sampling time $K$ is large enough, $\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is convergent by probability to the conditional expectation $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$, that is, for any positive $\epsilon_{2}$, we have,
@@
Let $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})=\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ and $\text{Var}(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t}))=\sigma$ for all $k$, the variance of $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})$ is as follows.
$\quad\text{Var}\left(\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})\right)=\text{Var}\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})\right)$ (23)
@@
The expectation of estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is convergent by probability to the expectation of true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for any positive $\epsilon_{2}$, we have,
@@
$\quad\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$ (26)
@@
$\quad\left.-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$
@@
$\quad\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\right|.$
@@
When the sampling number $K$ goes to infinity, the mean of the samplings converges to the true conditional expectation, for a small positive number $\epsilon_{2}$, we have,
$\small\mathop{\lim}_{k\rightarrow\infty}\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{2}\right)=1.$ (27)
The larger $K$ is, the closer the mean of samplings is to the true conditional expectation, which indicates the lower bound of the difference between the estimation conditional expectation and the true conditional function is zero.
V-B Error bound of approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$
@@
(Hoeffding’s inequality) [47] Let $u_{1}$,…$u_{K}$ be i.i.d. random variables such that $a\leq u_{k}\leq b$. Then for any $\epsilon_{3}>0$,
@@
$\quad\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}\left(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})]\right)\right|\geq\epsilon_{3}\right)$ (30)
@@
$\mathbf{x}^{(k)}_{\bar{S}}$ is sampled from $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the expectation of $f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is equal to $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ for any sample. Thus, Eq.(29) is obtained. ∎
The error bound of approximating $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ is discussed as follows.
@@
For any small positive $\epsilon_{2}$, the following inequality can be constructed,
$\small\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\mathbb{E}\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]\leq\epsilon_{2}.$ (31)
@@
$\quad\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{S}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})+f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right.\right.\right.$
$\quad\left.\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right.\right.$
$\quad\left.\left.+\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left(f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right)\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right.\right.$
$\quad\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$
@@
$\quad\left.+\left|o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$
$\leq\epsilon_{1}f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})+\epsilon_{1}^{\prime}$ (32)
$\leq\epsilon_{2}.$
According to Assumptions 1 and 2, $\mathbb{E}\left[\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}}\right]=0$ and $\mathbb{E}\left[o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\leq\epsilon_{1}^{\prime}$, we have,

### v4

@@ V Analysis on Error bound
In this section, the error bound of the contribution function in EmSHAP is discussed. Theorem 1 indicates that the expectation of the mean absolute deviation (MAD) between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ has an upper bound of $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$, where $K$ is the number of Monte Carlo sampling, $\epsilon_{2}$ is a small positive number related to the absolute probability density ratio between the estimated conditional density and the true data-generating density. For comparison, Theorems 2 and 3 present the error bounds of two competitive methods, i.e., KernelSHAP [14] and VAEAC [18].
@@
The parameters of the estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ are unbiased estimates of the parameters of the true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for a given positive $\epsilon_{1}$, the probability density ratio can be described as follows,
@@
Assumption 1 holds as EBM can be flexibly parameterized for arbitrary probability distributions [50, 51]. More specifically, the proposed GRU coupled energy-based model consists of an energy network and a GRU network. The energy network is an MLP with a skip connection and the GRU network is a type of gated recurrent network. According to Theorem 2.5 in Ref.[52], with weighted path norm uniformly bounded, by setting appropriate network width and depth, MLP with skip connection can approximate any function in Barron space [^2] Barron space is a function space that consists of functions which can be approximated by smooth combinations of simple basis functions. On the other hand, Theorem 3.2 in Ref.[53] shows the universal approximation ability of GRU.
Remark 1.
In practical applications, Assumption 1 may not hold strictly, particularly in high-dimensional or non-stationary data settings where conditional density estimation is inherently challenging. For high-dimensional case, increasing the amount of training data can alleviate this problem to some extent. For non-stationary data, Assumption 1 will be violated and the estimation accuracy will deteriorate if proper measures are not considered. A potential way to mitigate this scenario is to introduce standard preprocessing techniques such as differencing or detrending, another solution is to develop an adaptive version of EmSHAP. Furthermore, Assumption 3 enforces boundedness of $v(S)$ and $\hat{y}$ within $[0,1]$, which simplifies the theoretical error bound. In practice, this assumption can be easily achieved through normalization or scaling to ensure that the values remain within the specified range.
Before analyzing the upper error bound of EmSHAP, we first demonstrate the asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})]$, followed by the derivation of their error bounds.
V-A Asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})]$
@@
(Consistency) From the weak law of large numbers, when the sampling time $K$ is large enough, $\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is convergent by probability to the conditional expectation $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})]$, that is, for any positive $\epsilon_{2}$, we have,
@@
Let $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})=\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ and $\text{Var}(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t}))$ $\sigma$ for all $k$, the variance of $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})$ is as follows.
$\qquad\text{Var}\left(\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})\right)=\text{Var}\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})\right)$ (20)
@@
The expectation of estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is convergent by probability to the expectation of true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for a positive $\epsilon_{2}$ that is related to $\epsilon_{1}$ and $\epsilon_{1}^{\prime}$, we have,
@@
$\qquad\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$ (23)
@@
$\qquad\left.-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$
@@
$\qquad\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\right|.$
@@
When the sampling number $K$ goes to infinity, the mean of the samplings converges to the true conditional expectation, for a given positive number $\epsilon_{2}$, we have,
$\mathop{\lim}_{k\rightarrow\infty}\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{2}\right)$ (24)
$=1.$
The larger $K$ is, the closer the mean of the samplings is to the true conditional expectation, which indicates that the lower bound of the difference between the estimated conditional expectation and the true conditional function is zero.
V-B Error bound of approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})]$
@@
(Hoeffding’s inequality) [55] Let $u_{1}$$u_{K}$ be i.i.d. random variables such that $a\leq u_{k}\leq b$. Then for any $\epsilon_{3}>0$,
@@
$\qquad\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}\left(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})]\right)\right|\geq\epsilon_{3}\right)$ (27)
@@
$\mathbf{x}^{(k)}_{\bar{S}}$ is sampled from $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the expectation of $f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is equal to $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})]$ for any sample. Thus, Eq.(26) is obtained. ∎
The error bound of approximating $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})]$ is discussed as follows.
@@
For a given small positive $\epsilon_{2}$, the following inequality can be constructed,
$\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\mathbb{E}\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$ (28)
$\leq\epsilon_{2}.$
@@
$\qquad\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{S}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
[…]
$\qquad\left.\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
[…]
$\qquad\left.\left.+\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left(f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right)\right|\right]$
[…]
$\qquad\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$ (29)
@@
$\qquad\left.+\left|o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$
$\leq\epsilon_{1}f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})+\epsilon_{1}^{\prime}\leq\epsilon_{2}.$
According to Assumptions 1 and 2, $\mathbb{E}[\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}}]=0$ and $\mathbb{E}[o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})]\leq\epsilon_{1}^{\prime}$, we have,